# Scaling Up Liquid-Resistance Liquid-Capacitance Networks for Efficient Sequence Modeling

### v2

@@ Abstract
We present LrcSSM, a nonlinear recurrent model that processes long sequences as fast as today’s linear state-space layers. By forcing the state-transition matrix to be diagonal and learned at every step, the full sequence can be solved in parallel with a single prefix-scan, giving $\mathcal{O}(TD)$ time and memory and only $\mathcal{O}(\log T)$ sequential depth, for input-sequence length $T$ and a state dimension $D$. Moreover, LrcSSM offers a formal gradient-stability guarantee that other input-varying systems such as Liquid-S4 and Mamba do not provide. Lastly, for network depth $L$, as the forward and backward passes cost $\Theta(T\,D\,L)$ FLOPs, with its low sequential depth and parameter count $\Theta(D\,L)$, the model follows the compute-optimal scaling law regime ($\beta\approx 0.42$) recently observed for Mamba, outperforming quadratic-attention Transformers at equal compute while avoiding the memory overhead of FFT-based long convolutions. We show that on a series of long-range forecasting tasks, LrcSSM outperforms LRU, S5 and Mamba.
@@ 1 Introduction
With the advent of linear state space models (LSSMs), more and more architectures have emerged, with increasingly better accuracy and efficiency. While LSSMs can be efficiently parallelized, for example with the aid of the parallel scan operator, this is considerably more difficult for traditional, nonlinear state-space models (NSSMs). This lead to a decreasing interest in NSSMs, although these should arguably capture input correlations in a more refined way through their state.
Fortunately, recent work has shown how to apply the parallel scan operator to NSSMs, by linearizing them in every time step, and by implementing this idea in their DEER framework [24]. Unfortunately, the state-transition matrix (the Jacobian of the NSSM) was not diagonal, which precluded scaling it up to very long sequences. Subsequent work however, succeeded to scale up NSSMs by simply taking the diagonal of the Jacobian matrix, and stabilizing the DEER updates with trust regions. They called this method ELK (evaluating Levenberg-Marquardt via Kalman) [7].
In this paper, we propose an alternative approach to scaling up NSSMs. Instead of disregarding the non diagonal elements of the NSSM Jacobian, which might contain important information about the multistep interaction among neurons along feedback loops, we learn an NSSM whose state-transition matrix is constrained to be diagonal, and whose entries depend on both the current state and current input. As for LSSMs, our main intuition is that intricate loops of the NSSM state-transition (neural connectivity) matrix, can be well summarized by its complex eigenvalues. After all, the synaptic parameters define constant matrices that can be themselves diagonalized.
To test our idea, we modified and scaled up LRCs (liquid-resistance, liquid-capacitance neural networks), a bio-inspired NSSM [4] that considerably increased LTCs (liquid time constant networks) accuracy while also decreasing their convergence time  [13], by capturing saturation effects in biological neurons, and accounting for the state-and-input dependent nature of their capacitance. Most importantly, we introduce an inherent diagonalization to LRCs forcing the system’s Jacobian to be diagonal. This modification enables an exact ELK update - rather than an approximation.
Our experimental results on the Heartbeat, SelfRegulationSCP1, SelfRegulationSCP2, EthanolConcentration, MotorImagery, and EigenWorms long-sequence benchmarks, show that LrcSSMs are either on par or outperform state of the art LSSMs such as NRDE, NCDE, Log-NCDE, LRU, S5, S6, Mamba, LinOSS-IMEX, and LinOSS-IM. The EthanolConcentration benchmark in particular, seems to have the most intricate input correlations, a fact that will need future examination.
@@
  • To the best of our knowledge, we are the first to show how to scale up a bioinspired RNN to a competitive NSSM on long sequences and compare it to state of the art LSSMs.
  • We discuss in detail how to scale LRCs to LrcSSMs having a diagonal nonlinear state-and-input dependent state-transition matrix and inherently a diagonal Jacobian matrix.
  • We demonstrate that LrcSSMs can capture long-horizon tasks in a very competitive fashion on a set of standard benchmarks used to assess LSSMs accuracy and efficiency.
  • We show that LrcSSMs consistently outperform many of the state-of-the-art LSSMs, including LRU, S5, S6, and Mamba, especially on the EthanolConcentration benchmark.
Figure 1: Liquid-Resistance Liquid-Capacitance NSSM (LrcSSM) architecture. The input sequence of length $T$ and input dimension $p$ is first passed through an input encoder, followed by a normalization layer. The core component is a non-linear, state-and-input dependent LRC with hidden dimension $D$ and sequence length $T$. This NSSM is computed by a parallelizable iterative linearization method. The final state values are then processed by an MLP, with a skip connection added to preserve information flow. The LrcSSM block can be stacked and repeated an arbitrary number of times (we use ${2,4,6}$ layers in our experiments). A post-normalization layer is applied before the output is passed to the decoder, which produces the final output.
@@ 2 Background
Here we introduce the necessary background for understanding LrcSSMs: Firstly, the bioinspired nonlinear liquid networks LTCs, STCs, and LRCs, known for their dynamic expressivity. Secondly, the parallelization techniques enabling efficient training of traditionally sequential NSSMs.
@@ 2.1 Bio-inspired Liquid Neural Networks
Electrical Equivalent Circuits (EECs) are simplified models defining the dynamic behavior of the membrane potential (MP) of a postsynaptic neuron, as a function of the MP of its presynaptic neurons and external input signals [17, 36]. In ML, EECs with chemical synapses are termed liquid time constant networks (LTCs) [23, 13]. For a neuron $i$ with $m$ presynaptic neurons of MPs $x$ and $n$ inputs of value $u$, the forget conductance $f_{i}(x,u)$ and update conductance $z_{i}(x,u)$ are defined as:
${\color[rgb]{0,.5,.5}f_{i}(x,u)}$ $=\sum_{j=1}^{m+n}g_{ji}^{max}\sigma(a_{ji}y_{j}+b_{ji})+g_{i}^{leak}$ (1)
${\color[rgb]{0,0,1}z_{i}(x,u)}$ $=\sum_{j=1}^{m+n}k_{ji}^{max}\sigma(a_{ji}y_{j}+b_{ji})+g_{i}^{leak},$ (2)
@@
where $y\,{=}\,[x,u]$ concatenates the MP (state) of all neurons and the inputs. In Equation (1), $g_{ji}^{max}$ represents the maximum synaptic channel conductance, $a_{ji}$ and $b_{ji}$ parameterize the sigmoidal activation governing channel openness, and $g_{i}^{leak}$ is the leaking conductance. In Equation (2), $k_{ji}^{max}\,{=}\,g_{ji}^{max}e_{ji}^{rev}{/}e_{i}^{leak}$, where $e_{ji}^{rev}$ is the synaptic reversal potential (equilibrium membrane potential) and $e_{i}^{leak}$ is the leaking potential. Since $g_{ji}^{max}\geq 0$, the sign of $k_{ji}^{max}$ depends on $e_{ji}^{rev}{/}e_{i}^{leak}$.
@@
LTC-Equation (4) states that the rate of change of $x_{i}$ of neuron $i$, is the sum of its forget current $-f_{i}\,x_{i}$ and its update current $z_{i}\,e_{i}^{leak}$. LTCs ignore saturation aspects, which were introduced in saturated LTCs (STCs) in [3]. As ${\color[rgb]{0,.5,.5}f_{i}(x,u)}$ is positive, it is saturated with a sigmoid, and as ${\color[rgb]{0,0,1}z_{i}(x,u)}$ is either positive or negative, it is saturated with a tanh. Saturation is capturd in STC-Equation (5).
@@
Finally, both LTCs and STCs assume that the membrane capacitance is constant, and for simplicity equal to 1 in Equations (4)-(5), as the capacitance can be learned jointly with the other parameters. However, this assumption does not hold in biological neurons. In reality, the capacitance has a nonlinear dependence on the MP of presynaptic neurons and the external input as they both may cause the neuron to deform [16, 33, 21]. This behavior can be modeled by the following elastance:
$\sigma({\color[rgb]{1,.5,0}\epsilon_{i}(x,u)})=\sigma(\sum_{j=1}^{m+n}w_{ji}y_{j}+v_{j}),$ (3)
@@
where $y=[x,u]$, as before. LRCs incorporate this biological behavior of neurons, by introducing the elastance (which is the reciprocal of the capacitance) as a multiplicative term in LRC-Equation (6).
@@
[…]
The time constant of LRCs is $\frac{1}{RC}=\sigma({\color[rgb]{0,.5,.5}f_{i}(x,u)})\,\sigma({\color[rgb]{1,.5,0}\epsilon_{i}(x,u)})$, which factors into a liquid resistance $R=1/\sigma({\color[rgb]{0,.5,.5}f_{i}(x,u)})$ and a liquid capacitance $C=1/\sigma({\color[rgb]{1,.5,0}\epsilon_{i}(x,u)})$. While the resistive liquidity is the core of both LTCs and STCs, the capacitive liquidity acts as an additional control of the time constant.
@@
[…]
@@ 2.2 Parallelization Techniques
The DEER method [24] formulates next-state computation in NSSMs as a fixed-point problem and solves it using a parallel version of the Newton’s method. At each iteration step, DEER linearizes the NSSM. This approximation is widely effective across many domains, and often yields accurate estimates and fast convergence. The main limitation of DEER is the use of a square Jacobian, which does not scale up to long sequences when included in the parallel scan. The second limitation is its numerical instability, which arises from the nature of Newton’s method. In particular, the undamped version lacks global convergence guarantees and often diverges in practice [37, 7].
@@
To stabilize its convergence, Quasi-DEER leverages a connection between the Levenberg-Marquardt algorithm and Kalman smoothing in their ELK (Evaluating Levenberg-Marquardt with Kalman) algorithm [7]. This stabilization of the Newton iteration by constraining the step size within a trust region, prevents large and numerically unstable updates. As a result, updates are computed using a parallel Kalman smoother, with a running time that is logarithmic in the length of the sequence. Algorithm 1 below, presents this method [7].
Algorithm 1 ELK
[…]
@@
[…]
8:         $J_{s}\leftarrow\textsc{Diag}(J_{s})$
[…]
@@ 3 Scaling Up Non-linear LRCs
A scalable DEER or ELK approximation, first computes the dense Jacobian of the NSSM, as shown in Line 7 of Algorithm 1, and then extracts its diagonal as shown in Line 8. This results in a quasi approximation of the original DEER technique, called Quasi-DEER and Quasi-ELK [7].
@@
Instead of following this approach, we directly modify the underlying nonlinear LRC of Equation (6), such that its Jacobian is diagonal by the formulation itself. The main idea of this modification is that the state-connectivity submatrices $a^{x}$, $w^{x}$, $g^{max,x}$, and $k^{max,x}$ of the state-and-input-connectivity matrices $a$, $w$, $g^{max}$, and $k^{max}$, are constant parameter matrices that are theselves diagonlizable. Consequently, all cross terms are zeroed out in the LRC through diagonalization. Accordingly, we learn the complex diagonal matrices directly, instead.
As a result, our own algorithm is no longer a quasi-approximation, as we do not explicitly remove nondiagonal entries. Instead we learn their contribution to the dynamics, within the complex eigenvalues of the diagonal. Consequently, Line 8, $J_{s}\,{\leftarrow}\,\text{Diag}(J_{s})$ of Algorithm 1, is not needed anymore, and the update computations become more efficient. In this way, we retain the best of both approaches: A much more precise, more stable, and more scalable, parallelization technique.
@@ 3.1 Proposed Model
In order to achieve a diagonal Jacobian for the LRCs by default, we first modify the Equations (1)-(3), by splitting their summation terms into a state-dependent and an input-dependent group, respectively. For the former, we only keep the self-loop synaptic parameters, and zero out all the cross-state synaptic parameters in the associated matrices. For the latter we keep the influence of all external inputs $u$ through their cross-input synaptic parameters, as this part is zeroed out anyway in the Jacobian. To highlight the separation of the terms, we include an extra superscript $x$ for the learnable parameters in the state-dependent part, and the superscript of $u$ for the parameters in the input-dependent part. This separation results in Equations (8)-(10).
As a consequence, instead of keeping cross-synaptic activations, where each individual synapse between neuron $j$ and $i$ has its own $g_{ji}^{max}$, $b_{ji}$ and $k_{ji}^{max}$ as it was in Equation (1) and (2), we now only keep the self-loop neural activations, where the synaptic parameters from the same neuron are equal. Note that instead of the ij indices, we have only the $j$ index in Equations (8) and (9).
We denote the modified equations of the LRCs with an asterisk. This gives us the following equations for the ${\color[rgb]{0,.5,.5}f_{i}^{*}(x_{i},u)}$, ${\color[rgb]{0,0,1}z_{i}^{*}(x_{i},u)}$, and ${\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},u)}$ terms:
${\color[rgb]{0,.5,.5}f_{i}^{*}(x_{i},u)}$ $=\underbrace{g_{i}^{max,x}\sigma(a_{i}^{x}x_{i}+b_{i}^{x})}_{\text{$x_{i}$ state-dependent}}$ $+\underbrace{g_{i}^{max,u}\sigma(\sum_{j=1}^{n}a_{ji}^{u}u_{j}+b_{j}^{u})}_{\text{$u$ input-dependent}}+g_{i}^{leak}$ (8)
${\color[rgb]{0,0,1}z_{i}^{*}(x_{i},u)}$ $=\underbrace{k_{i}^{max,x}\sigma(a_{i}^{x}x_{i}+b_{i}^{x})}_{\text{$x_{i}$ state-dependent}}$ $+\underbrace{k_{i}^{max,u}\sigma(\sum_{j=1}^{n}a_{ji}^{u}u_{j}+b_{j}^{u})}_{\text{$u$ input-dependent}}+g_{i}^{leak}$ (9)
${\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},u)}$ $=\underbrace{w_{i}^{x}x_{i}+v_{i}^{x}}_{\text{$x_{i}$ state-dependent}}$ $+\underbrace{\sum_{j=1}^{n}w_{ji}^{u}u_{j}+v_{j}^{u}}_{\text{$u$ input-dependent}}$ (10)
[…]
@@
[…]
@@
$\mathbf{A}(\mathbf{x},\mathbf{u})=\text{diag}\begin{bmatrix}-\sigma({\color[rgb]{0,.5,.5}f_{1}^{*}(x_{1},u)})\sigma({\color[rgb]{1,.5,0}\epsilon_{1}^{*}(x_{1},u)})\\ ...\\ -\sigma({\color[rgb]{0,.5,.5}f_{i}^{*}(x_{i},u)})\sigma({\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},u)})\\ ...\\ -\sigma({\color[rgb]{0,.5,.5}f_{m}^{*}(x_{m},u)})\sigma({\color[rgb]{1,.5,0}\epsilon_{m}^{*}(x_{m},u)})\\ \end{bmatrix},$ (13)
@@
$\mathbf{b}(\mathbf{x},\mathbf{u})=\begin{bmatrix}\tau({\color[rgb]{0,0,1}z_{1}^{*}(x_{1},u)})\sigma({\color[rgb]{1,.5,0}\epsilon_{1}^{*}(x_{1},u)})\,e_{1}^{leak}\\ ...\\ \tau({\color[rgb]{0,0,1}z_{i}^{*}(x_{i},u)})\sigma({\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},u)})\,e_{i}^{leak}\\ ...\\ \tau({\color[rgb]{0,0,1}z_{m}^{*}(x_{m},u)})\sigma({\color[rgb]{1,.5,0}\epsilon_{m}^{*}(x_{m},u)})\,e_{m}^{leak}\\ \end{bmatrix}.$ (14)
[…]

### v6

# Parallelization of Non-linear State-Space Models: Scaling Up Liquid-Resistance Liquid-Capacitance Networks for Efficient Sequence Modeling

@@ Abstract
We present LrcSSM, a non-linear recurrent model that processes long sequences as fast as today’s linear state-space layers. By forcing its Jacobian matrix to be diagonal, the full sequence can be solved in parallel, giving $\mathcal{O}(TD)$ computational work and memory and only $\mathcal{O}(\log T)$ sequential depth, for input-sequence length $T$ and a state dimension $D$. Moreover, LrcSSM offers a formal gradient-stability guarantee that other input-varying systems such as Liquid-S4 and Mamba do not provide. Importantly, the diagonal Jacobian structure of our model results in no performance loss compared to the original model with dense Jacobian, and the approach can be generalized to other non-linear recurrent models, demonstrating broader applicability. On a suite of long-range forecasting tasks, we demonstrate that LrcSSM outperforms Transformers, LRU, S5, and Mamba.
@@ 1 Introduction
With the advent of linear structured state space models (SSMs), more and more architectures have emerged, with increasingly better accuracy and efficiency. While they can be efficiently parallelized, for example with the aid of the parallel scan operator, this is considerably more difficult for traditional, non-linear models. This led to a decreasing interest in non-linear recurrent neural networks (RNNs), although these should arguably capture input correlations in a more refined way through their state.
Fortunately, recent work has shown how to apply the parallel scan operator to non-linear RNNs, by linearizing them in every time step, and by implementing this idea in their DEER framework [27]. Unfortunately, the state-transition matrix (the Jacobian of the model) was not diagonal, which precluded scaling it up to very long sequences. Subsequent work however, succeeded to scale up non-linear RNNs by simply taking the diagonal of the Jacobian matrix, and stabilizing the DEER updates with trust regions. They called this method ELK (evaluating Levenberg-Marquardt via Kalman) [8].
In this paper, we propose an alternative approach to scaling up non-linear RNNs. Instead of disregarding the non-diagonal elements of the model’s Jacobian, which might contain important information about the multistep interaction among neurons along feedback loops, we learn a model whose Jacobian matrix is constrained to be diagonal, and whose entries depend on both the current state and current input. As for linear SSMs, our main intuition is that intricate loops of the non-linear SSM state (neural connectivity) matrix can be well summarized by its complex eigenvalues. After all, the synaptic parameters define constant matrices that can be themselves diagonalized.
To test our idea, we modified and scaled up LRCs (liquid-resistance, liquid-capacitance neural networks), a bio-inspired non-linear RNN [5] that considerably increased LTCs (liquid time-constant networks) accuracy while also decreasing their convergence time [14], by capturing saturation effects in biological neurons, and accounting for the state-and-input dependent nature of their capacitance. Most importantly, we introduce an inherent diagonal form to LRCs, forcing the system’s Jacobian to be diagonal. This modification enables exact updates during parallelization - rather than approximations.
Our experimental results on the Heartbeat, SelfRegulationSCP1, SelfRegulationSCP2, EthanolConcentration, MotorImagery, and EigenWorms long-sequence benchmarks show that LrcSSMs are either on par or outperform state-of-the-art SSMs such as NRDE, NCDE, Log-NCDE, LRU, S5, S6, Mamba, LinOSS-IMEX and LinOSS-IM and Transformer variants.
@@
  • We discuss in detail how to scale LRCs to LrcSSMs having a diagonal non-linear state-and-input dependent state matrix, resulting in an inherently diagonal Jacobian matrix, which allows exact computations via efficient parallelization.
  • We demonstrate that LrcSSMs can capture long-horizon tasks in a very competitive fashion on a set of standard benchmarks used to assess accuracy and efficiency of SSMs.
  • We show that LrcSSMs consistently outperform many of the state-of-the-art SSMs, including LRU, S5, S6, and Mamba, especially on the EthanolConcentration benchmark.
  • We show how our diagonal model approach can be generalized to other non-linear recurrent models, broadening the impact of its design.
footnotetext: Our code is available at https://github.com/MoniFarsang/LrcSSM
Figure 1: Liquid-Resistance Liquid-Capacitance SSM (LrcSSM) architecture. The input sequence of length $T$ and input dimension $p$ is first passed through an input encoder, followed by a normalization layer. The core component is a non-linear, state-and-input dependent LRC with hidden dimension $D$ and sequence length $T$. This SSM is computed by a parallelizable iterative linearization method. The final state values are then processed by an MLP, with a skip connection added to preserve information flow. The LrcSSM block can be stacked and repeated an arbitrary number of times (we use 2,4,6{2,4,6} layers in our experiments). A post-normalization layer is applied before the output is passed to the decoder, which produces the final output.
@@ 2 Background
Here we introduce the necessary background for understanding LrcSSMs: Firstly, the bio-inspired non-linear liquid networks - LTCs, STCs, and LRCs - known for their dynamic expressivity. Secondly, the parallelization techniques enabling efficient training of traditionally sequential non-linear models.
@@ 2.1 Bio-inspired Liquid Neural Networks
Electrical Equivalent Circuits (EECs) are simplified models defining the dynamic behavior of the membrane potential (MP) of a postsynaptic neuron, as a function of the MP of its presynaptic neurons and external input signals [19, 42]. In ML, EECs with chemical synapses are termed liquid time-constant networks (LTCs) [26, 14]. For a neuron $i$ with $m$ presynaptic neurons of MPs $\mathbf{x}$ and $n$ inputs of value $\mathbf{u}$, the forget conductance $f_{i}(\mathbf{x},\mathbf{u})$ and update conductance $z_{i}(\mathbf{x},\mathbf{u})$ are defined as:
${\color[rgb]{0,.5,.5}f_{i}(\mathbf{x},\mathbf{u})}$ $=\sum_{j=1}^{m+n}g_{ji}^{max}\sigma(a_{ji}y_{j}+b_{ji})+g_{i}^{leak}$ (1)
${\color[rgb]{0,0,1}z_{i}(\mathbf{x},\mathbf{u})}$ $=\sum_{j=1}^{m+n}k_{ji}^{max}\sigma(a_{ji}y_{j}+b_{ji})+g_{i}^{leak},$ (2)
@@
where $\mathbf{y}\,{=}\,[\mathbf{x},\mathbf{u}]$ concatenates the MP (state) of all neurons and the inputs. In Equation (1), $g_{ji}^{max}$ represents the maximum synaptic channel conductance, $a_{ji}$ and $b_{ji}$ parameterize the sigmoidal activation governing channel openness, and $g_{i}^{leak}$ is the leaking conductance. In Equation (2), $k_{ji}^{max}\,{=}\,g_{ji}^{max}e_{ji}^{rev}{/}e_{i}^{leak}$, where $e_{ji}^{rev}$ is the synaptic reversal potential (equilibrium membrane potential) and $e_{i}^{leak}$ is the leaking potential. Since $g_{ji}^{max}\geq 0$, the sign of $k_{ji}^{max}$ depends on $e_{ji}^{rev}{/}e_{i}^{leak}$.
@@
LTC-Equation (4) states that the rate of change of $x_{i}$ of neuron $i$, is the sum of its forget current $f_{i}\,x_{i}$ and its update current $z_{i}\,e_{i}^{leak}$. LTCs ignore saturation aspects, which were introduced in saturated LTCs (STCs) in [4]. As $\color[rgb]{0,.5,.5}f_{i}(\mathbf{x},\mathbf{u})$} is positive, it is saturated with a sigmoid, and as $\color[rgb]{0,0,1}z_{i}(\mathbf{x},\mathbf{u})$} is either positive or negative, it is saturated with a tanh. Saturation is captured in STC-Equation (5).
@@
Finally, both LTCs and STCs assume that the membrane capacitance is constant, and for simplicity equal to 1 in Equations (4)-(5), as the capacitance can be learned jointly with the other parameters. However, this assumption does not hold in biological neurons. In reality, the capacitance has a non-linear dependence on the MP of presynaptic neurons and the external input as they both may cause the neuron to deform [18, 38, 23]. This behavior can be modeled by the following elastance:
$\sigma({\color[rgb]{1,.5,0}\epsilon_{i}(\mathbf{x},\mathbf{u})})=\sigma(\sum_{j=1}^{m+n}w_{ji}y_{j}+v_{j}),$ (3)
@@
where $\mathbf{y}=[\mathbf{x},\mathbf{u}]$, as before. LRCs incorporate this biological behavior of neurons, by introducing the elastance (which is the reciprocal of the capacitance) as a multiplicative term in LRC-Equation (6).
$\text{LTC:\quad}\dot{x}_{i}$ $=\hskip 14.22636pt-{\color[rgb]{0,.5,.5}f_{i}(\mathbf{x},\mathbf{u})}\,x_{i}\hskip 4.2679pt+\hskip 8.53581pt{\color[rgb]{0,0,1}z_{i}(\mathbf{x},\mathbf{u})}\,e_{i}^{leak}$ (4)
$\text{STC:\quad}\dot{x}_{i}$ $=\hskip 4.2679pt-\sigma({\color[rgb]{0,.5,.5}f_{i}(\mathbf{x},\mathbf{u})})\,x_{i}+\tau({\color[rgb]{0,0,1}z_{i}(\mathbf{x},\mathbf{u})})\,e_{i}^{leak}$ (5)
$\text{LRC:\quad}\dot{x}_{i}$ $=(-\sigma({\color[rgb]{0,.5,.5}f_{i}(\mathbf{x},\mathbf{u})})\,x_{i}+\tau({\color[rgb]{0,0,1}z_{i}(\mathbf{x},\mathbf{u})})\,e_{i}^{leak})\,\sigma({\color[rgb]{1,.5,0}\epsilon_{i}(\mathbf{x},\mathbf{u})})$ (6)
@@
$\text{LRC:\quad}\mathbf{x}_{t}=\mathbf{x}_{t-1}+\Delta t\,\mathbf{\dot{x}}_{t-1}$ (7)
@@ 2.2 Parallelization Techniques
The DEER method [27] formulates next-state computation in non-linear RNNs as a fixed-point problem and solves it using a parallel version of the Newton’s method. At each iteration step, DEER linearizes the model. This approximation is widely effective across many domains, and often yields accurate estimates and fast convergence. The main limitation of DEER is the use of a square Jacobian, which does not scale up to long sequences when included in the parallel scan. The second limitation is its numerical instability, which arises from the nature of Newton’s method. In particular, the undamped version lacks global convergence guarantees and often diverges in practice [43, 8].
@@
To stabilize its convergence, the connection between the Levenberg-Marquardt algorithm and Kalman smoothing is leveraged in the ELK (Evaluating Levenberg-Marquardt with Kalman) algorithm [8]. This stabilisation of the Newton iteration by constraining the step size within a trust region prevents large and numerically unstable updates. As a result, updates are computed using a parallel Kalman smoother, with a running time that is logarithmic in the length of the sequence. Algorithm 1 below, presents these methods [8].
Algorithm 1 DEER/ELK method with optional quasi approximation [8]
1:procedure ParallelizeRNN($f,s_{0},\text{init\_guess},\text{tol},\text{method},\text{quasi}$)
@@
3:  $\text{states}\leftarrow\text{init\_guess}$
4:  while $\text{diff}>\text{tol}$ do
5:   $\text{shifted\_states}\leftarrow[s_{0},\text{states}[:-1]]$
6:   $f_{s}\leftarrow f(\text{shifted\_states})$
7:   $J_{s}\leftarrow\textsc{GetJacobians}(f,\text{shifted\_states})$
8:   $\text{{if} quasi: }J_{s}\leftarrow\textsc{Diag}(J_{s})$
9:   $b_{s}\leftarrow f_{s}-J_{s}\cdot\text{shifted\_states}$
10:   $\text{new\_states}\leftarrow\text{method}(J_{s},b_{s},\text{states},s_{0})$
11:   $\text{diff}\leftarrow\lVert\,\text{states}-\text{new\_states}\rVert_{\infty}$
12:   $\text{states}\leftarrow\text{new\_states}$
@@ 3 Scaling Up Non-linear LRCs
A scalable DEER or ELK approximation, first computes the dense Jacobian of the non-linear RNN, as shown in Line 7 of Algorithm 1, and then extracts its diagonal as shown in Line 8. This results in a quasi approximation of the original DEER/ELK technique, called quasi-DEER and quasi-ELK [8].
@@
Instead of following this approach, we directly modify the underlying non-linear LRC-Equation (6), such that its Jacobian is diagonal by the model formulation itself. The main idea of this modification is that the state-connectivity submatrices $a^{x}$, $w^{x}$ and $g^{max,x}$ are constant parameter matrices that are themselves diagonalizable. Consequently, all cross terms are zeroed out in the LRC through diagonalization. Accordingly, we learn the complex diagonal matrices (vectors) directly, instead.
As a result, our own algorithm is no longer a quasi-approximation, as we do not explicitly remove non-diagonal entries of the Jacobian during parallelization. Instead, we inherently learn their contribution to the dynamics in the model, within the complex eigenvalues of the diagonal. Consequently, Line 8, $J_{s}\,{\leftarrow}\,\text{Diag}(J_{s})$ of Algorithm 1, is not needed anymore, and the update computations become more efficient. In this way, we retain the best of both approaches: A much more precise, more stable, and more scalable, parallelization technique by model design.
@@ 3.1 Proposed Model
In order to achieve a diagonal Jacobian for the LRCs by model design, we first modify the Equations (1)-(3), by splitting their summation terms into a state-dependent and an input-dependent group, respectively. For the former, we only keep the self-loop synaptic parameters, and zero out all the cross-state synaptic parameters in the associated matrices. For the latter we keep the influence of all external inputs $u$ through their cross-input synaptic parameters, as this part is zeroed out anyway in the Jacobian. To highlight the separation of the terms, we include an extra superscript $x$ for the learnable parameters in the state-dependent part, and the superscript of $u$ for the parameters in the input-dependent part. This separation results in Equations (8)-(10).
As a consequence, instead of keeping cross-synaptic activations, where each individual synapse between neuron $j$ and $i$ has its own $g_{ji}^{max}$, $b_{ji}$ and $k_{ji}^{max}$ as it was in Equation (1) and (2), we now only keep the self-loop neural activations, where the synaptic parameters from the same neuron are equal. Note that instead of the ij indices, we have only the $i$ index for $g^{max}$ and $k^{max}$, and $j$ for $b$ in Equations (8) and (9).
We denote the modified equations of the LRCs with an asterisk. This gives us the following equations for the $\color[rgb]{0,.5,.5}f_{i}^{*}(x_{i},\mathbf{u})$}, $\color[rgb]{0,0,1}z_{i}^{*}(x_{i},\mathbf{u})$}, and $\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},\mathbf{u})$} terms:
${\color[rgb]{0,.5,.5}f_{i}^{*}(x_{i},\mathbf{u})}$ $=\underbrace{g_{i}^{max,x}\sigma(a_{i}^{x}x_{i}+b_{i}^{x})}_{\text{$x_{i}$ state-dependent}}$ $+\underbrace{g_{i}^{max,u}\sigma(\sum_{j=1}^{n}a_{ji}^{u}u_{j}+b_{j}^{u})}_{\text{$\mathbf{u}$ input-dependent}}+g_{i}^{leak}$ (8)
${\color[rgb]{0,0,1}z_{i}^{*}(x_{i},\mathbf{u})}$ $=\underbrace{k_{i}^{max,x}\sigma(a_{i}^{x}x_{i}+b_{i}^{x})}_{\text{$x_{i}$ state-dependent}}$ $+\underbrace{k_{i}^{max,u}\sigma(\sum_{j=1}^{n}a_{ji}^{u}u_{j}+b_{j}^{u})}_{\text{$\mathbf{u}$ input-dependent}}+g_{i}^{leak}$ (9)
${\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},\mathbf{u})}$ $=\underbrace{w_{i}^{x}x_{i}+v_{i}^{x}}_{\text{$x_{i}$ state-dependent}}$ $+\underbrace{\sum_{j=1}^{n}w_{ji}^{u}u_{j}+v_{j}^{u}}_{\text{$\mathbf{u}$ input-dependent}}$ (10)
$\text{LrcSSM: \quad}\dot{x}_{i}$ $=-\sigma({\color[rgb]{0,.5,.5}f_{i}^{*}(x_{i},\mathbf{u})})\sigma({\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},\mathbf{u})})\,x_{i}+\tau({\color[rgb]{0,0,1}z_{i}^{*}(x_{i},\mathbf{u})})\sigma({\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},\mathbf{u})})\,e_{i}^{leak}$ (11)
@@
$\text{LrcSSM: \quad}\dot{\mathbf{x}}=\mathbf{A}(\mathbf{x},\mathbf{u})\mathbf{x}+\mathbf{b}(\mathbf{x},\mathbf{u}),$ (12)
@@
$\mathbf{A}(\mathbf{x},\mathbf{u})=\text{diag}\begin{bmatrix}-\sigma({\color[rgb]{0,.5,.5}f_{1}^{*}(x_{1},\mathbf{u})})\sigma({\color[rgb]{1,.5,0}\epsilon_{1}^{*}(x_{1},\mathbf{u})})\\ ...\\ -\sigma({\color[rgb]{0,.5,.5}f_{i}^{*}(x_{i},\mathbf{u})})\sigma({\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},\mathbf{u})})\\ ...\\ -\sigma({\color[rgb]{0,.5,.5}f_{m}^{*}(x_{m},\mathbf{u})})\sigma({\color[rgb]{1,.5,0}\epsilon_{m}^{*}(x_{m},\mathbf{u})})\\ \end{bmatrix}$, (13)
@@
$\mathbf{b}(\mathbf{x},\mathbf{u})=\begin{bmatrix}\tau({\color[rgb]{0,0,1}z_{1}^{*}(x_{1},\mathbf{u})})\sigma({\color[rgb]{1,.5,0}\epsilon_{1}^{*}(x_{1},\mathbf{u})})\,e_{1}^{leak}\\ ...\\ \tau({\color[rgb]{0,0,1}z_{i}^{*}(x_{i},\mathbf{u})})\sigma({\color[rgb]{1,.5,0}\epsilon_{i}^{*}(x_{i},\mathbf{u})})\,e_{i}^{leak}\\ ...\\ \tau({\color[rgb]{0,0,1}z_{m}^{*}(x_{m},\mathbf{u})})\sigma({\color[rgb]{1,.5,0}\epsilon_{m}^{*}(x_{m},\mathbf{u})})\,e_{m}^{leak}\\ \end{bmatrix}$. (14)
This reduced diagonal $\textbf{A}(\textbf{x},\textbf{u})$ form of Equation (13) and the reduced version of $\textbf{b}(\textbf{x},\textbf{u})$ in Equation (14) results in a diagonal Jacobian matrix which makes the parallelizable iterative state updates exact and efficient, that is, this is not anymore a quasi-approximation of the Jacobian.